# Probabilistic Robustness for Free? Revisiting Training via a Benchmark

### v1

@@ 4.1 Empirical Analysis of Evaluation Results
RT methods underperform AT methods on PR performance. While RT methods effectively improve PR compared with ERM, they still lag behind AT methods, as shown in Table 3. Moreover, their PR performance drops sharply as the perturbation radius increases, cf. Fig. 2 (c) (d), showing limited generalization beyond the norm-ball setting in training. Notably, our analysis of the CVaR codebase from Robey et al. (2022) suggests a potential evaluation inconsistency: it includes misclassified original clean samples when computing $\textit{ProbAcc}(\rho)$ rather than focusing on the robustness of correctly classified original samples. We have corrected this in our experiments. This view is also supported by prior studies, e.g., Li et al. (2024), which states: “When considering the adversarial robustness of a wrongly classified sample $\bm{x}$, the robustness should be 0”. Similarly, Chen & Lee (2024) emphasizes that “AEs must satisfy the constraints: the original examples are classified correctly while the predictions of the AEs are wrong”. In other words, if a sample is misclassified by the model initially, its ability to resist perturbations and maintain the wrong prediction (i.e., being “robustly wrong”) does not constitute meaningful robustness.
@@
Training efficiency varies widely. Table 3 shows that Corruption method is highly efficient, generating only a single random sample per training example, whereas CVaR is computationally expensive due to the need for multiple samples to approximate the perturbation distribution. Standard AT methods exhibit intermediate and more consistent training times. AT-PR is the most costly, as it generates a set of PGD-based AEs and selects the one maximizing the all-AE region, leading to a substantial increase in training time (as noted by the complexity analysis in Zhang et al. (2025)). To further investigate efficiency-oriented approaches, we additionally integrate the filtering-based method of Chen & Lee (2024) into PRBench as a case study (Table 11). This aligns with ongoing work on efficient robustness, and our extensible design enables inclusion of future methods.
@@ A.1 Experiment Setup
For the MNIST dataset, we employ a four-layer CNN, consisting of two convolutional layers followed by two fully connected layers, as described in Robey et al. (2022). For the CIFAR-10, CIFAR-100, CINIC-10 and SVHN datasets, we independently train VGG-19, ResNet-18, and WideResNet-28-10 on each dataset, and additionally include Vision Transformer (ViT) and Data-efficient Image Transformer (Deit) for CIFAR-10. For the TinyImageNet dataset, we train ResNet-18 and ResNet-34 using the same training configuration as for CIFAR-10. For the ImageNet-50 dataset, we train a ResNet-18 model separately. All models are trained using stochastic gradient descent (SGD) with a momentum of 0.9 and a weight decay coefficient of $3.5\times 10^{-3}$. Training is performed for 100 epochs (200 for ImageNet-50) with an initial learning rate of 0.01, which is decayed by a factor of 10 at epochs 75 and 90, following the setup in Wang et al. (2019); Robey et al. (2022).
@@ A.2 Training Algorithms Hyperparameter Setting
  • TRADES. We used the same step size and number of steps as described above for PGD. Additionally, we applied a weight of $\lambda=6.0$ for all datasets, following the approach in Zhang et al. (2019); Robey et al. (2022).
  • MART. We used the same step size and number of steps as described above for PGD. Additionally, we applied a weight of $\lambda=5.0$ for all datasets, following the approach in Wang et al. (2019).
  • ALP. Follow the original work Kannan et al. (2018), we set $\lambda=1$ for all datasets, except $\lambda=0.01$ for SVHN, MNIST, ImageNet-50 and TinyImageNet.
@@
  • CVaR. We set the number of perturbed samples drawn from the uniform distribution to 20, following the original setup in Robey et al. (2022).
@@
  • AT-PR. The number of PGD-based AE candidates is set to 5 while other parameters follow the original setup in Zhang et al. (2025).
  • Chen & Lee (2024). Following the standard efficient parameter configuration, we integrate it into TRADES with $m=1\text{–}8$ and $k=10$ (10-step attack).
@@ A.4 Model and Dataset Selection
In $\mathtt{PRBench}$, we consider three major model architectures: (1) plain CNNs, such as VGG, representing early feedforward convolutional networks; (2) residual CNNs, exemplified by ResNet, which introduce skip connections to enable deeper architectures; and (3) transformer-based models, including Vision Transformer (ViT) Dosovitskiy et al. (2021) and Data-efficient Image Transformer (Deit) Touvron et al. (2021), which leverage self-attention mechanisms for image representation learning. In our experiments, we conduct different training methods (See Table. 1) across a diverse set of models: VGG-19, SimpleCNN (4-layer CNN), ResNet-18, ResNet-34, Wide-ResNet-28-10, ViT, and DeiT, covering various scales such as ViT-Small, ViT-Base, ViT-Large, DeiT-Tiny, and DeiT-Small. See Table 5 for details.
@@
  • MNIST Deng (2012) is a classic dataset of handwritten digits, consisting of 60,000 training images and 10,000 test images. Each image is a 28x28 pixel grayscale image, and the dataset is categorized into 10 classes (0-9).
  • SVHN Netzer et al. (2011) (Street View House Numbers) is a real-world dataset derived from Google Street View house numbers. It contains 73,257 training and 26,032 test images, each sized 32×32 and labeled into 10 digit classes. Compared to MNIST, SVHN is more challenging due to background clutter and variation, making it suitable for evaluating robustness under real-world conditions.
  • CIFAR-10 Krizhevsky et al. (2009) is a widely used image classification dataset with 60,000 32x32 color images across 10 classes. The dataset is split into 50,000 training images and 10,000 test images. Compared to MNIST and SVHN, CIFAR-10 presents greater variability in content and background, making it a standard benchmark for evaluating generalization across diverse object classes.
  • CIFAR-100 Krizhevsky et al. (2009) extends CIFAR-10 by increasing the number of classes to 100, with each class containing 600 images. The dataset is also composed of 60,000 images (50,000 for training and 10,000 for testing) with 32x32 size. Its finer-grained classification task poses a greater challenge, requiring models to generalize across a broader and more diverse set of object categories.
  • CINIC-10 Darlow et al. (2018) is derived from CIFAR-10 and downsampled ImageNet, with all images resized to 32×32 resolution. It contains 270,000 images evenly split into training, validation, and test sets (90,000 each). Like CIFAR-10, it has 10 classes, but the larger scale and more diverse image sources make it a more challenging and realistic benchmark.
  • TinyImageNet Le & Yang (2015) is a subset of the ImageNet dataset, containing 100,000 training, 10,000 validation, and 10,000 test images, all resized to 64×64 pixels and spanning 200 object classes. With higher resolution and a larger number of categories, TinyImageNet poses a more complex classification challenge, making it well-suited for evaluating model scalability and robustness in realistic settings.
  • ImageNet-50 Deng et al. (2009) is a 50-class subset of ImageNet for efficiency. It consists of approximately 65,000 training images and 2,500 validation images at 224×224 resolution. Despite being smaller than full ImageNet, ImageNet-50 retains considerable diversity and complexity, serving as a scalable benchmark for robustness evaluation.
@@ A.5 Related Benchmarks
Several robustness platforms have been developed to support AR evaluation by implementing popular attack methods, such as FoolBox Rauber et al. (2017), AdverTorch Ding et al. (2019), Cleverhan Papernot et al. (2016a), AdvBox Goodman et al. (2020), ART Nicolae et al. (2018), SecML Melis et al. (2019), DeepRobust Li et al. (2020), etc. In addition, some AR benchmarks have also been established  Ling et al. (2019); Guo et al. (2023); Croce et al. (2021); Tang et al. (2021); Dong et al. (2020); Liu et al. (2025). Despite these efforts, existing work remains confined to AR evaluation, overlooking the more practical perspective of PR. Consequently, they fail to reveal the inherent relationship between AR and PR, thereby hindering the establishment of a holistic and unified understanding of robustness progress. Compared to the aforementioned benchmarks, $\mathtt{PRBench}$ provides a more comprehensive framework in several key aspects: (1) it is the first benchmark to emphasize the PR, providing a formalized definition of PR along with a summary of commonly used evaluation metrics and associated training methods; (2) it includes an extensive evaluation of widely used models and datasets, covering representative training methods designed for both AR and PR, enabling direct comparisons under PR metrics and revealing the performance of these methods in both adversarial and probabilistic scenarios; and (3) it presents in-depth analyses of the generalization error for PR from empirical and theoretical perspectives, which, to the best of our knowledge, has not been explored in prior benchmarks. These insights help improve DL model robustness in practical applications by offering a more complete view from both adversarial and probabilistic perspectives.
@@ Appendix B Supplementary Experiments
Table 11: Evaluation of an efficient AT method Chen & Lee (2024) using ResNet-18 on CIFAR-10 & CIFAR-100. The method employs a simple yet effective data filtering mechanism to enhance training efficiency and robustness. Integrated into TRADES with m=0m=088 and $k=10$ (10-step attack), it reduces computational cost, lowers AR/PR generalization error, and improves PR performance while maintaining comparable clean accuracy and AR. This experiment also illustrates the extensibility of PRBench, which is designed to incorporate and compare future efficient AT methods.
@@ Appendix C PR-targeted training method
Some recent works have focused on designing training techniques specifically to improve PR. Wang et al. (2021) proposed a simple augmentation-based approach referred to as corruption training, where the model is trained on randomly perturbed examples. This process is equivalent to standard AT training, except that each input $\bm{x}_{i}$ is perturbed by a sample drawn from a uniform distribution within an $\ell_{\infty}$ norm ball of radius $\gamma$, rather than using worst-case AEs. They show that corruption training improves PR and reduces generalization error under PR evaluation settings. The corresponding optimization objective is formalized as:
@@
Later, Robey et al. (2022) introduced an optimization objective based on the Conditional Value-at-Risk (CVaR) to improve PR. Specifically, for a loss function $\mathcal{L}$ and a continuous distribution $Pr$, CVaR can be interpreted as the expected value of $\mathcal{L}$ over the tail of the distribution. It admits the following convex, variational characterization, which allows it to be optimized via stochastic gradient-based methods. Based on this, they formalized the training objective as:
@@
Building upon CVaR, Zhang et al. (2024a) propose Probabilistic Risk-averse Robust Learning with Stochastic Search (PRASS), which trains a risk-averse model by mitigating the Entropic Value-at-Risk (EVaR) over the captured worst-case perturbation distribution. Unlike CVaR, which focuses solely on the tail samples, EVaR further capitalizes on the entire distribution of samples. Specifically, EVaR with risk level $\rho\in(0,1]$, denoted as $\text{EVaR}_{\rho}(\mathcal{L}(\bm{\delta}))$, is defined as the infimum over $\alpha>0$ of the Chernoff bound for $\mathcal{L}(\bm{\delta})$ with respect to the random variable $\bm{\delta}$. They then proposed a risk-averse robust learning paradigm to optimize the upper bound, and formalized the training objective as:
@@
The latest PR-targeted work Zhang et al. (2025), though aimed at PR, is not based on the perturbation risk function in Def. 2. Instead, it is closer to the traditional AT formulation in Eq. 3, introducing a new min–max objective designed to identify AEs that lie in the largest all-AE region $k$ (i.e., $\textit{PR}(\bm{x}+\bm{\delta},k)=0$). They achieve this by designing a numerical algorithm that first uses PGD to generate a diverse set of AEs at different local optima, i.e., $\mathcal{S}=\{\bm{x}^{\prime}_{1},\bm{x}^{\prime}_{2},\dots,\bm{x}^{\prime}_{m}\}$. For each $\bm{x}^{\prime}_{i}$, gradient descent is then performed toward the nearest decision boundary, and the AE with the largest distance to its nearest boundary is selected, corresponding to the largest all-AE region. Formally, for a model $f_{\bm{\theta}}$ trained on dataset $\mathcal{D}$, the objective of AT-PR is:
@@ D.1 AR Evaluation Metrics
AR is a critical aspect of model robustness and has been extensively studied in previous benchmark works. To measure AR, adversarial attacks (Eq. 2) are typically adopted to generate AEs $\bm{x}^{\prime}$ for each input $\bm{x}$ in the test dataset. In $\mathtt{PRBench}$, we consider four white-box attacks: two variants of Projected Gradient Descent attack with 10 and 20 iterations ($\text{PGD}^{10}$ and $\text{PGD}^{20}$), the C&W attack Carlini & Wagner (2017), and Auto-attack Croce & Hein (2020). Formally, AR is computed as the classification accuracy of the target model on adversarially perturbed inputs across the test dataset:
@@ D.2 PR Evaluation Metrics
PR was first introduced by Webb et al. (2019), which proposed a formal framework to quantify PR under random perturbations (Def. 1). Computing the exact PR requires taking an expectation over the perturbation distribution, which is typically intractable. Therefore, a Monte Carlo approximation is adopted by uniformly sampling a large number of perturbations within a norm ball of radius $\gamma$. The target model then predicts the labels of these perturbed inputs to determine whether they are AEs $\bm{x}^{\prime}$. PR is thus defined as the proportion of non-adversarial samples among all sampled perturbations. Formally, given a test dataset $\mathcal{D}$ and a perturbation budget $\gamma$, PR is estimated as:
@@
Later, Robey et al. (2022) proposed quantile accuracy $\textit{ProbAcc}(\rho)$ to evaluate PR for individual inputs using the same perturbation strategy (uniform sampling around each input). Unlike $\textit{PR}_{\mathcal{D}}(\gamma)$, which estimates the mean PR over all correctly classified clean test samples, this metric focuses on a given robustness tolerance threshold $\rho$. Specifically, $\textit{ProbAcc}(\rho)$ measures the proportion of inputs whose $\textit{PR}(\bm{x},\gamma)$ exceeds the threshold $1-\rho$, formally defined as follows:

### v2

@@ 4.1 Empirical Analysis of Evaluation Results
RT methods underperform AT methods on PR performance. While RT methods effectively improve PR compared with ERM, they still lag behind AT methods, as shown in Table 3. Moreover, their PR performance drops sharply as the perturbation radius increases, cf. Fig. 2 (c) (d), showing limited generalization beyond the norm-ball setting in training. Notably, our analysis of the CVaR codebase from robey2022probabilistically suggests a potential evaluation inconsistency: it includes misclassified original clean samples when computing $\textit{ProbAcc}(\rho)$ rather than focusing on the robustness of correctly classified original samples. We have corrected this in our experiments. This view is also supported by prior studies, e.g., li2024great, which states: “When considering the adversarial robustness of a wrongly classified sample $\bm{x}$, the robustness should be 0”. Similarly, chen2024data emphasizes that “AEs must satisfy the constraints: the original examples are classified correctly while the predictions of the AEs are wrong”. In other words, if a sample is misclassified by the model initially, its ability to resist perturbations and maintain the wrong prediction (i.e., being “robustly wrong”) does not constitute meaningful robustness.
@@
Training efficiency varies widely. Table 3 shows that Corruption method is highly efficient, generating only a single random sample per training example, whereas CVaR is computationally expensive due to the need for multiple samples to approximate the perturbation distribution. Standard AT methods exhibit intermediate and more consistent training times. AT-PR is the most costly, as it generates a set of PGD-based AEs and selects the one maximizing the all-AE region, leading to a substantial increase in training time (as noted by the complexity analysis in zhang2025adversarial). To further investigate efficiency-oriented approaches, we additionally integrate the filtering-based method of chen2024data into PRBench as a case study (Table 11). This aligns with ongoing work on efficient robustness, and our extensible design enables inclusion of future methods.
@@ A.1 Experiment Setup
For the MNIST dataset, we employ a four-layer CNN, consisting of two convolutional layers followed by two fully connected layers, as described in robey2022probabilistically. For the CIFAR-10, CIFAR-100, CINIC-10 and SVHN datasets, we independently train VGG-19, ResNet-18, and WideResNet-28-10 on each dataset, and additionally include Vision Transformer (ViT) and Data-efficient Image Transformer (Deit) for CIFAR-10. For the TinyImageNet dataset, we train ResNet-18 and ResNet-34 using the same training configuration as for CIFAR-10. For the ImageNet-50 dataset, we train a ResNet-18 model separately. All models are trained using stochastic gradient descent (SGD) with a momentum of 0.9 and a weight decay coefficient of $3.5\times 10^{-3}$. Training is performed for 100 epochs (200 for ImageNet-50) with an initial learning rate of 0.01, which is decayed by a factor of 10 at epochs 75 and 90, following the setup in wang2019improving; robey2022probabilistically.
@@ A.2 Training Algorithms Hyperparameter Setting
  • TRADES. We used the same step size and number of steps as described above for PGD. Additionally, we applied a weight of $\lambda=6.0$ for all datasets, following the approach in zhang2019theoretically; robey2022probabilistically.
  • MART. We used the same step size and number of steps as described above for PGD. Additionally, we applied a weight of $\lambda=5.0$ for all datasets, following the approach in wang2019improving.
  • ALP. Follow the original work kannan2018adversarial, we set $\lambda=1$ for all datasets, except $\lambda=0.01$ for SVHN, MNIST, ImageNet-50 and TinyImageNet.
@@
  • CVaR. We set the number of perturbed samples drawn from the uniform distribution to 20, following the original setup in robey2022probabilistically.
@@
  • AT-PR. The number of PGD-based AE candidates is set to 5 while other parameters follow the original setup in zhang2025adversarial.
  • chen2024data. Following the standard efficient parameter configuration, we integrate it into TRADES with $m=1\text{–}8$ and $k=10$ (10-step attack).
@@ A.4 Model and Dataset Selection
In $\mathtt{PRBench}$, we consider three major model architectures: (1) plain CNNs, such as VGG, representing early feedforward convolutional networks; (2) residual CNNs, exemplified by ResNet, which introduce skip connections to enable deeper architectures; and (3) transformer-based models, including Vision Transformer (ViT) dosovitskiy2021an and Data-efficient Image Transformer (Deit) touvron2021training, which leverage self-attention mechanisms for image representation learning. In our experiments, we conduct different training methods (See Table. 1) across a diverse set of models: VGG-19, SimpleCNN (4-layer CNN), ResNet-18, ResNet-34, Wide-ResNet-28-10, ViT, and DeiT, covering various scales such as ViT-Small, ViT-Base, ViT-Large, DeiT-Tiny, and DeiT-Small. See Table 5 for details.
@@
  • MNIST deng2012mnist is a classic dataset of handwritten digits, consisting of 60,000 training images and 10,000 test images. Each image is a 28x28 pixel grayscale image, and the dataset is categorized into 10 classes (0-9).
  • SVHN netzer2011reading (Street View House Numbers) is a real-world dataset derived from Google Street View house numbers. It contains 73,257 training and 26,032 test images, each sized 32×32 and labeled into 10 digit classes. Compared to MNIST, SVHN is more challenging due to background clutter and variation, making it suitable for evaluating robustness under real-world conditions.
  • CIFAR-10 krizhevsky2009learning is a widely used image classification dataset with 60,000 32x32 color images across 10 classes. The dataset is split into 50,000 training images and 10,000 test images. Compared to MNIST and SVHN, CIFAR-10 presents greater variability in content and background, making it a standard benchmark for evaluating generalization across diverse object classes.
  • CIFAR-100 krizhevsky2009learning extends CIFAR-10 by increasing the number of classes to 100, with each class containing 600 images. The dataset is also composed of 60,000 images (50,000 for training and 10,000 for testing) with 32x32 size. Its finer-grained classification task poses a greater challenge, requiring models to generalize across a broader and more diverse set of object categories.
  • CINIC-10 darlow2018cinic is derived from CIFAR-10 and downsampled ImageNet, with all images resized to 32×32 resolution. It contains 270,000 images evenly split into training, validation, and test sets (90,000 each). Like CIFAR-10, it has 10 classes, but the larger scale and more diverse image sources make it a more challenging and realistic benchmark.
  • TinyImageNet le2015tiny is a subset of the ImageNet dataset, containing 100,000 training, 10,000 validation, and 10,000 test images, all resized to 64×64 pixels and spanning 200 object classes. With higher resolution and a larger number of categories, TinyImageNet poses a more complex classification challenge, making it well-suited for evaluating model scalability and robustness in realistic settings.
  • ImageNet-50 deng2009imagenet is a 50-class subset of ImageNet for efficiency. It consists of approximately 65,000 training images and 2,500 validation images at 224×224 resolution. Despite being smaller than full ImageNet, ImageNet-50 retains considerable diversity and complexity, serving as a scalable benchmark for robustness evaluation.
@@ A.5 Related Benchmarks
Several robustness platforms have been developed to support AR evaluation by implementing popular attack methods, such as FoolBox rauber2017foolbox, AdverTorch ding2019advertorch, Cleverhan papernot2016technical, AdvBox goodman2020advbox, ART nicolae2018adversarial, SecML melis2019secml, DeepRobust li2020deeprobust, etc. In addition, some AR benchmarks have also been established  ling2019deepsec; guo2023comprehensive; croce2020robustbench; tang2021robustart; dong2020benchmarking; liu2025comprehensive. Despite these efforts, existing work remains confined to AR evaluation, overlooking the more practical perspective of PR. Consequently, they fail to reveal the inherent relationship between AR and PR, thereby hindering the establishment of a holistic and unified understanding of robustness progress. Compared to the aforementioned benchmarks, $\mathtt{PRBench}$ provides a more comprehensive framework in several key aspects: (1) it is the first benchmark to emphasize the PR, providing a formalized definition of PR along with a summary of commonly used evaluation metrics and associated training methods; (2) it includes an extensive evaluation of widely used models and datasets, covering representative training methods designed for both AR and PR, enabling direct comparisons under PR metrics and revealing the performance of these methods in both adversarial and probabilistic scenarios; and (3) it presents in-depth analyses of the generalization error for PR from empirical and theoretical perspectives, which, to the best of our knowledge, has not been explored in prior benchmarks. These insights help improve DL model robustness in practical applications by offering a more complete view from both adversarial and probabilistic perspectives.
@@ Appendix B Supplementary Experiments
Table 11: Evaluation of an efficient AT method chen2024data using ResNet-18 on CIFAR-10 & CIFAR-100. The method employs a simple yet effective data filtering mechanism to enhance training efficiency and robustness. Integrated into TRADES with m=0m=088 and $k=10$ (10-step attack), it reduces computational cost, lowers AR/PR generalization error, and improves PR performance while maintaining comparable clean accuracy and AR. This experiment also illustrates the extensibility of PRBench, which is designed to incorporate and compare future efficient AT methods.
@@ Appendix C PR-targeted training method
Some recent works have focused on designing training techniques specifically to improve PR. wang2021statistically proposed a simple augmentation-based approach referred to as corruption training, where the model is trained on randomly perturbed examples. This process is equivalent to standard AT training, except that each input $\bm{x}_{i}$ is perturbed by a sample drawn from a uniform distribution within an $\ell_{\infty}$ norm ball of radius $\gamma$, rather than using worst-case AEs. They show that corruption training improves PR and reduces generalization error under PR evaluation settings. The corresponding optimization objective is formalized as:
@@
Later, robey2022probabilistically introduced an optimization objective based on the Conditional Value-at-Risk (CVaR) to improve PR. Specifically, for a loss function $\mathcal{L}$ and a continuous distribution $Pr$, CVaR can be interpreted as the expected value of $\mathcal{L}$ over the tail of the distribution. It admits the following convex, variational characterization, which allows it to be optimized via stochastic gradient-based methods. Based on this, they formalized the training objective as:
@@
Building upon CVaR, zhang2024prass propose Probabilistic Risk-averse Robust Learning with Stochastic Search (PRASS), which trains a risk-averse model by mitigating the Entropic Value-at-Risk (EVaR) over the captured worst-case perturbation distribution. Unlike CVaR, which focuses solely on the tail samples, EVaR further capitalizes on the entire distribution of samples. Specifically, EVaR with risk level $\rho\in(0,1]$, denoted as $\text{EVaR}_{\rho}(\mathcal{L}(\bm{\delta}))$, is defined as the infimum over $\alpha>0$ of the Chernoff bound for $\mathcal{L}(\bm{\delta})$ with respect to the random variable $\bm{\delta}$. They then proposed a risk-averse robust learning paradigm to optimize the upper bound, and formalized the training objective as:
@@
The latest PR-targeted work zhang2025adversarial, though aimed at PR, is not based on the perturbation risk function in Def. 2. Instead, it is closer to the traditional AT formulation in Eq. 3, introducing a new min–max objective designed to identify AEs that lie in the largest all-AE region $k$ (i.e., $\textit{PR}(\bm{x}+\bm{\delta},k)=0$). They achieve this by designing a numerical algorithm that first uses PGD to generate a diverse set of AEs at different local optima, i.e., $\mathcal{S}=\{\bm{x}^{\prime}_{1},\bm{x}^{\prime}_{2},\dots,\bm{x}^{\prime}_{m}\}$. For each $\bm{x}^{\prime}_{i}$, gradient descent is then performed toward the nearest decision boundary, and the AE with the largest distance to its nearest boundary is selected, corresponding to the largest all-AE region. Formally, for a model $f_{\bm{\theta}}$ trained on dataset $\mathcal{D}$, the objective of AT-PR is:
@@ D.1 AR Evaluation Metrics
AR is a critical aspect of model robustness and has been extensively studied in previous benchmark works. To measure AR, adversarial attacks (Eq. 2) are typically adopted to generate AEs $\bm{x}^{\prime}$ for each input $\bm{x}$ in the test dataset. In $\mathtt{PRBench}$, we consider four white-box attacks: two variants of Projected Gradient Descent attack with 10 and 20 iterations ($\text{PGD}^{10}$ and $\text{PGD}^{20}$), the C&W attack carlini2017towards, and Auto-attack croce2020reliable. Formally, AR is computed as the classification accuracy of the target model on adversarially perturbed inputs across the test dataset:
@@ D.2 PR Evaluation Metrics
PR was first introduced by webb2018statistical, which proposed a formal framework to quantify PR under random perturbations (Def. 1). Computing the exact PR requires taking an expectation over the perturbation distribution, which is typically intractable. Therefore, a Monte Carlo approximation is adopted by uniformly sampling a large number of perturbations within a norm ball of radius $\gamma$. The target model then predicts the labels of these perturbed inputs to determine whether they are AEs $\bm{x}^{\prime}$. PR is thus defined as the proportion of non-adversarial samples among all sampled perturbations. Formally, given a test dataset $\mathcal{D}$ and a perturbation budget $\gamma$, PR is estimated as:
@@
Later, robey2022probabilistically proposed quantile accuracy $\textit{ProbAcc}(\rho)$ to evaluate PR for individual inputs using the same perturbation strategy (uniform sampling around each input). Unlike $\textit{PR}_{\mathcal{D}}(\gamma)$, which estimates the mean PR over all correctly classified clean test samples, this metric focuses on a given robustness tolerance threshold $\rho$. Specifically, $\textit{ProbAcc}(\rho)$ measures the proportion of inputs whose $\textit{PR}(\bm{x},\gamma)$ exceeds the threshold $1-\rho$, formally defined as follows: